# April 3, 2025

### v1

@@ 4. Existence of solutions for Problem $\mathcal{P}(\Omega)$
By the uniform positive-definiteness of the elasticity tensors $\{A^{ijk\ell}\}$ and $\{B^{ijk\ell}\}$, Korn’s inequality (Theorem 2.1), (4), and (5), the following estimate holds for $m$ sufficiently large:
(6) $\rho\|\dot{\bm{u}}^{m}_{\kappa}(t)\|_{\bm{L}^{2}(\Omega)}^{2}+\dfrac{\|\bm{u}^{m}_{\kappa}(t)\|_{\bm{V}(\Omega)}^{2}}{2c_{0}C_{e}^{(1)}}+\dfrac{1}{2c_{0}C_{e}^{(2)}}\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{\chi_{(0,T_{0})}(t)}{2\kappa}\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}$
$\leq\rho\|\bm{u}_{1}\|_{\bm{L}^{2}(\Omega)}^{2}+\dfrac{c_{0}C_{e}^{(2)}}{2}\|\bm{f}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0}^{m})e_{i\|j}(\bm{u}_{0}^{m})\,\mathrm{d}x$
$\quad+\rho\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2c_{0}C_{e}^{(1)}}\int_{0}^{t}\|\bm{u}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau$
$\quad+\dfrac{1}{2\kappa}\int_{0}^{t}\chi_{(0,T_{0})}(\tau)\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(\tau)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}\,\mathrm{d}\tau.$
@@
$y(t):=\rho\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2c_{0}C_{e}^{(1)}}\int_{0}^{t}\|\bm{u}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2\kappa}\int_{0}^{t}\chi_{(0,T_{0})}(\tau)\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(\tau)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}\,\mathrm{d}\tau,$
@@
$b\equiv\max\left\{1,\rho,\dfrac{c_{0}C_{e}^{(2)}}{2}\right\}\left(\|\bm{u}_{1}\|_{\bm{L}^{2}(\Omega)}^{2}+\|\bm{f}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x\right)\geq 0,$
gives the following uniform upper bound
(7) $\rho\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2c_{0}C_{e}^{(1)}}\int_{0}^{t}\|\bm{u}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau$
$\quad+\dfrac{1}{2\kappa}\int_{0}^{t}\chi_{(0,T_{0})}(\tau)\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(\tau)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}\,\mathrm{d}\tau$
$\leq T\max\left\{1,\rho,\dfrac{c_{0}C_{e}^{(2)}}{2}\right\}\left(\|\bm{u}_{1}\|_{\bm{L}^{2}(\Omega)}^{2}+\|\bm{f}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x\right)e^{T},$
for all $t\in[0,T]$. Therefore, we obtain that
@@
and, moreover, by (6), there exists a positive uniform constant $L$ such that
(9) $0\leq\left\|\chi_{(0,T_{0})}\left\{[\bm{I}+u^{m}_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(0,T;L^{2}(\Gamma))}\leq\sqrt{\kappa Tbe^{T}}.$
@@
$|\langle\ddot{\bm{u}}^{m}_{\kappa}(t),\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}|\leq\|\bm{f}(t)\|_{\bm{L}^{2}(\Omega)}+\left(\max_{\overline{\Omega}}|A^{ijk\ell}|\right)\|\bm{u}^{m}_{\kappa}\|_{L^{\infty}(0,T;\bm{V}(\Omega))}$
$\quad+\left(\max_{\overline{\Omega}}|B^{ijk\ell}|\right)\|\dot{\bm{u}}^{m}_{\kappa}(t)\|_{\bm{V}(\Omega)}+\dfrac{1}{\kappa}\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\bm{q}\right\|_{\bm{L}^{2}(\Gamma)},$
@@
(10) $\|\ddot{\bm{u}}^{m}_{\kappa}\|_{L^{2}(0,T;\bm{V}^{\ast}(\Omega))}\leq C\left(1+\dfrac{1}{\kappa}\right).$
In light of the assumed admissibility of the applied body force $\bm{f}$ (Definition 3.1), we have that for a.a. $T_{0}<t<T$, there exists a constant $C=C(T_{0})>0$, independent of $m$, $t$ and $\kappa$, such that:
(11) $\|\ddot{\bm{u}}^{m}_{\kappa}\|_{L^{2}(T_{0},T;\bm{V}^{\ast}(\Omega))}\leq C.$
@@
[…]
@@
An application of Theorem 8.28 of [29] to the fourth convergence of the process (12) gives:
(14) $\kappa^{-1}\chi_{(0,T_{0})}\left\{[\bm{I}+u^{m}_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\rightharpoonup\chi_{\kappa},\quad\textup{ in }L^{2}((0,T)\times\Omega)\textup{ as }m\to\infty.$
Thanks to Lemma 4.1, the first convergence of (13) and the weak convergence (14), Theorem 8.28 of [29], Theorem 8.62 of [29] and the monotonicity of the operator $\tilde{\mathcal{N}}$ defined in (1), we are in a position to apply Theorem 9.13-2 of [10] and, so, to obtain:
(15) $\chi_{\kappa}=\kappa^{-1}\chi_{(0,T_{0})}\left\{[\bm{I}+u_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\in L^{2}((0,T)\times\Omega).$
@@
$\quad-\dfrac{1}{\kappa}\int_{0}^{T}\chi_{(0,T_{0})}(t)\int_{\Omega}\left(\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)(v_{i}\bm{e}^{i}\cdot\bm{q})\,\mathrm{d}x\psi(t)\,\mathrm{d}t$
@@
for all $\bm{v}\in\bm{E}^{\tilde{m}}$. Thanks to Lemma 4.1, the convergence process (12), (15), the arbitrariness of $\psi\in\mathcal{D}(0,T)$, as well as the fact that
@@
a passage to the limit as $m\to\infty$ in (16) shows that $\bm{u}_{\kappa}$ is a solution to the penalised variational equations in Problem $\mathcal{P}_{\kappa}(\Omega)$.
@@
$\bm{u}_{\kappa}(0)=\bm{u}_{0},\quad\textup{ in }\bm{L}^{2}(\Omega).$
@@
The next step consists in showing that the sequence of solutions $\{\bm{u}_{\kappa}\}_{\kappa>0}$ of Problem $\mathcal{P}_{\kappa}(\Omega)$ converges in some suitable sense to a solution of Problem $\mathcal{P}(\Omega)$. Before establishing this result, that constitutes the main contribution of this paper, we need to establish some preparatory lemmas.
@@
The operator $\mathcal{A}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ defined by
$\langle\mathcal{A}\bm{w},\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}:=\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{w})e_{i\|j}(\bm{v})\,\mathrm{d}x,$
is linear and continuous. Moreover, also the corresponding time-dependent version $\tilde{\mathcal{A}}:L^{\infty}(0,T;\bm{V}(\Omega))\to L^{\infty}(0,T;\bm{V}^{\ast}(\Omega))$ defined pointwise by
$(\tilde{\mathcal{A}}\bm{w})(t):=\mathcal{A}(\bm{w}(t)),\quad\textup{ for all }\bm{w}\in L^{\infty}(0,T;\bm{V}(\Omega)),$
for a.a. $t\in(0,T)$ is linear and continuous.
Proof.
Lemma 4.3.
The operator $\mathcal{B}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ defined by
$\langle\mathcal{B}\bm{w},\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}:=\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{w})e_{i\|j}(\bm{v})\,\mathrm{d}x,$
is linear and continuous. Moreover, also the corresponding time-dependent version ~$\tilde{\mathcal{B}}:L^{2}(0,T;\bm{V}(\Omega))\to L^{2}(0,T;\bm{V}^{\ast}(\Omega))$ defined pointwise by
(~$(\tilde{\mathcal{B}}\bm{w})(t):=\mathcal{B}(\bm{w}(t)),\quad\textup{ for all }\bm{w}\in L^{2}(0,T;\bm{V}(\Omega)),$
for a.a. $t\in(0,T)$ is linear and continuous. ∎
Lemma 4.4.
Consider the sequence $\{\bm{u}_{\kappa}\}_{\kappa>0}$, where each $\bm{u}_{\kappa}$ is a solution of Problem $\mathcal{P}_{\kappa}(\Omega)$. Up to passing to a subsequence, the following convergences hold:
[…]
$\dot{\bm{u}}_{\kappa}$ $\rightharpoonup\dot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{V}(\Omega))\textup{ as }\kappa\to 0,$
$\ddot{\bm{u}}_{\kappa}$ $\rightharpoonup\ddot{\bm{u}},\quad\textup{ in }L^{2}(T_{0},T;\bm{V}^{\ast}(\Omega))\textup{ as }\kappa\to 0,$
$\chi_{(0,T_{0})}\{[\bm{I}+u_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\}^{-}$ $\to\chi_{(0,T_{0})}\{[\bm{I}+u_{i}\bm{e}^{i}]\cdot\bm{q}\}=0,\quad\textup{ in }L^{2}(0,T;\bm{L}^{2}(\Gamma))\textup{ as }\kappa\to 0,$
$\bm{u}_{\kappa}$ $\rightharpoonup\bm{u},\quad\textup{ in }\mathcal{C}^{0}([0,T];\bm{V}(\Omega))\textup{ as }\kappa\to 0,$
$\dot{\bm{u}}_{\kappa}(T)$ $\rightharpoonup\dot{\bm{u}}(T),\quad\textup{ in }\bm{V}^{\ast}(\Omega)\textup{ as }\kappa\to 0.$

### v3

@@ 4. Existence of solutions for Problem $\mathcal{P}(\Omega)$
By the uniform positive-definiteness of the elasticity tensors $\{A^{ijk\ell}\}_{i,j,k,\ell}$ and $\{B^{ijk\ell}\}_{i,j,k,\ell}$, Korn’s inequality (Theorem 2.1), (10), and (11), the following estimate holds for $m$ sufficiently large:
(12) $\rho\|\dot{\bm{u}}^{m}_{\kappa}(t)\|_{\bm{L}^{2}(\Omega)}^{2}+\dfrac{\|\bm{u}^{m}_{\kappa}(t)\|_{\bm{V}(\Omega)}^{2}}{2c_{0}C_{e}^{(1)}}+\dfrac{1}{2c_{0}C_{e}^{(2)}}\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2\kappa}\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}$
$\leq\rho\|\bm{u}_{1}^{m}\|_{\bm{L}^{2}(\Omega)}^{2}+\dfrac{c_{0}C_{e}^{(2)}}{2}\|\bm{f}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}+\dfrac{1}{2}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0}^{m})e_{i\|j}(\bm{u}_{0}^{m})\,\mathrm{d}x$
$\quad+\rho\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2c_{0}C_{e}^{(1)}}\int_{0}^{t}\|\bm{u}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2\kappa}\int_{0}^{t}\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(\tau)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}\,\mathrm{d}\tau.$
@@
$y(t):=\rho\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2c_{0}C_{e}^{(1)}}\int_{0}^{t}\|\bm{u}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2\kappa}\int_{0}^{t}\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(\tau)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}\,\mathrm{d}\tau,$
@@
$b\equiv\max\left\{\dfrac{1}{2},\rho,\dfrac{c_{0}C_{e}^{(2)}}{2}\right\}\left(\|\bm{u}_{1}^{m}\|_{\bm{L}^{2}(\Omega)}^{2}+\|\bm{f}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0}^{m})e_{i\|j}(\bm{u}_{0}^{m})\,\mathrm{d}x\right)\geq 0,$
gives the following upper bound
(13) $\rho\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2c_{0}C_{e}^{(1)}}\int_{0}^{t}\|\bm{u}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2\kappa}\int_{0}^{t}\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(\tau)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}\,\mathrm{d}\tau$
$\leq t\max\left\{\dfrac{1}{2},\rho,\dfrac{c_{0}C_{e}^{(2)}}{2}\right\}\left(\|\bm{u}_{1}^{m}\|_{\bm{L}^{2}(\Omega)}^{2}+\|\bm{f}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0}^{m})e_{i\|j}(\bm{u}_{0}^{m})\,\mathrm{d}x\right)e^{t},$
which can be easily made independent of $t$. Observe that the previous right hand side is bounded independently of $m$ in light of Theorem 4.13-1 in [15]. Therefore, combining (12) with (13) gives
@@
and, moreover, by (13) and Gronwall’s inequality (cf., e.g., [28] or Appendix B.2(j) in [21]), it results:
(15) $\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(0,T;L^{2}(\Gamma))}\leq\sqrt{\kappa Tbe^{T}}.$
@@
$2\rho|\langle\ddot{\bm{u}}^{m}_{\kappa}(t),\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}|\leq\|\bm{f}(t)\|_{\bm{L}^{2}(\Omega)}+\left(\max_{i,j,k,\ell}\|A^{ijk\ell}\|_{\mathcal{C}^{0}(\overline{\Omega})}\right)\|\bm{u}^{m}_{\kappa}(t)\|_{\bm{V}(\Omega)}$
$\quad+\left(\max_{i,j,k,\ell}\|B^{ijk\ell}\|_{\mathcal{C}^{0}(\overline{\Omega})}\right)\|\dot{\bm{u}}^{m}_{\kappa}(t)\|_{\bm{V}(\Omega)}+\dfrac{1}{\kappa}\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\bm{q}\right\|_{\bm{L}^{2}(\Gamma)},$
@@
(16) $\|\ddot{\bm{u}}^{m}_{\kappa}\|_{L^{2}(0,T;\bm{V}^{\ast}(\Omega))}\leq\dfrac{C}{2\rho}\left(1+\dfrac{1}{\sqrt{\kappa}}\right).$
@@
(17) $\bm{u}^{m}_{\kappa}\overset{\ast}{\rightharpoonup}\bm{u}_{\kappa},$ $\textup{ in }L^{\infty}(0,T;\bm{V}(\Omega))\textup{ as }m\to\infty,$
$\dot{\bm{u}}^{m}_{\kappa}\overset{\ast}{\rightharpoonup}\dot{\bm{u}}_{\kappa},$ $\textup{ in }L^{\infty}(0,T;\bm{L}^{2}(\Omega))\textup{ as }m\to\infty,$
$\dot{\bm{u}}^{m}_{\kappa}\rightharpoonup\dot{\bm{u}}_{\kappa},$ $\textup{ in }L^{2}(0,T;\bm{V}(\Omega))\textup{ as }m\to\infty,$
$\ddot{\bm{u}}^{m}_{\kappa}\rightharpoonup\ddot{\bm{u}}_{\kappa},$ $\textup{ in }L^{2}(0,T;\bm{V}^{\ast}(\Omega))\textup{ as }m\to\infty,$
$\kappa^{-1}\left\{[\bm{I}+u^{m}_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\rightharpoonup\chi_{\kappa},$ $\textup{ in }L^{2}(0,T;L^{2}(\Gamma))\textup{ as }m\to\infty.$
@@
An application of Theorem 8.28 of [37] to the fifth convergence of the process (17) gives:
(19) $\kappa^{-1}\left\{[\bm{I}+u^{m}_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\rightharpoonup\chi_{\kappa},\quad\textup{ in }L^{2}((0,T)\times\Gamma)\textup{ as }m\to\infty.$
Combining (19) with Lemma 4.2 gives:
(20) $\dfrac{1}{\kappa}\int_{0}^{T}\langle\mathcal{N}\bm{u}^{m}_{\kappa}(t),\bm{u}^{m}_{\kappa}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t$
$=-\dfrac{1}{\kappa}\int_{0}^{T}\int_{\Gamma}\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}(\bm{I}+u^{m}_{i,\kappa}\bm{e}^{i})\cdot\bm{q}\,\mathrm{d}\Gamma\,\mathrm{d}t$
$\quad+\dfrac{1}{\kappa}\int_{0}^{T}\int_{\Gamma}\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}(\bm{I}\cdot\bm{q})\,\mathrm{d}\Gamma\,\mathrm{d}t$
$\to-\int_{0}^{T}\int_{\Gamma}\chi_{\kappa}(t)(u_{i,\kappa}(t)\bm{e}^{i}\cdot\bm{q})\,\mathrm{d}\Gamma\,\mathrm{d}t,\quad\textup{ as }m\to\infty.$
Thanks to Lemma 4.1, the first convergence of (18) and the weak convergence (19), Theorem 8.28 of [37], Theorem 8.62 of [37] and the monotonicity of the operator $\tilde{\mathcal{N}}$ defined in (4) and (20), we are in a position to apply Theorem 12.5-2 of [15] and, so, to obtain:
(21) $\chi_{\kappa}=\kappa^{-1}\left\{[\bm{I}+u_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\in L^{2}((0,T)\times\Gamma).$
@@
$\quad-\dfrac{1}{\kappa}\int_{0}^{T}\int_{\Gamma}\left(\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)(v_{i}\bm{e}^{i}\cdot\bm{q})\,\mathrm{d}\Gamma\psi(t)\,\mathrm{d}t$
@@
for all $\bm{v}\in\bm{E}^{\tilde{m}}$. Thanks to Lemma 4.1, the convergence process (17), (21), the arbitrariness of $\psi\in\mathcal{D}(0,T)$, the fact that
@@
and the continuity of the operators $\mathcal{A}$ (Lemma 4.3) and $\mathcal{B}$ (Lemma 4.4), we obtain that a passage to the limit as $m\to\infty$ in (22) gives that $\bm{u}_{\kappa}$ is a solution to the penalised variational equations in Problem $\mathcal{P}_{\kappa}(\Omega)$.
@@
$\bm{u}_{\kappa}(0)=\bm{u}_{0},\quad\textup{ in }\bm{V}(\Omega).$
@@
The next step consists in showing that the family of solutions $\{\bm{u}_{\kappa}\}_{\kappa>0}$ of Problem $\mathcal{P}_{\kappa}(\Omega)$ admits a subsequence $\{\bm{u}_{\kappa_{n}}\}_{n=1}^{\infty}$ (where $\{\kappa_{n}\}_{n=1}^{\infty}$ denotes a sequence of positive real numbers that converges to zero as $n\to\infty$) that converges in some suitable sense to a solution of Problem $\mathcal{P}(\Omega)$.
Thanks to the energy estimates recovered in the proof of Theorem 4.3, we can extract further compactness, as asserted in the following lemma.
@@
Consider the sequence $\{\bm{u}_{\kappa}\}_{\kappa>0}$, where each $\bm{u}_{\kappa}$ is a solution of Problem $\mathcal{P}_{\kappa}(\Omega)$. Up to passing to a subsequence $\{\bm{u}_{\kappa_{n}}\}_{n=1}^{\infty}$ where $\{\kappa_{n}\}_{n=1}^{\infty}$ is such that $\kappa_{n}\to 0^{+}$ as $n\to\infty$, the following convergences hold:
(23) $\bm{u}_{\kappa_{n}}$ $\overset{\ast}{\rightharpoonup}\bm{u},\quad\textup{ in }L^{\infty}(0,T;\bm{V}(\Omega))\textup{ as }n\to\infty,$
$\dot{\bm{u}}_{\kappa_{n}}$ $\overset{\ast}{\rightharpoonup}\dot{\bm{u}},\quad\textup{ in }L^{\infty}(0,T;\bm{L}^{2}(\Omega))\textup{ as }n\to\infty,$
$\dot{\bm{u}}_{\kappa_{n}}$ $\rightharpoonup\dot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{V}(\Omega))\textup{ as }n\to\infty,$
$\{[\bm{I}+u_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\}^{-}$ $\to\{[\bm{I}+u_{i}\bm{e}^{i}]\cdot\bm{q}\}^{-}=0,\quad\textup{ in }L^{2}(0,T;\bm{L}^{2}(\Gamma))\textup{ as }n\to\infty,$
$\bm{u}_{\kappa_{n}}$ $\rightharpoonup\bm{u},\quad\textup{ in }\mathcal{C}^{0}([0,T];\bm{V}(\Omega))\textup{ as }n\to\infty.$